OpenStudy (anonymous):

Define the inverse cotangent function by restricting the domain of the cotangent function to the interval (0, π). Use a calculator to approximate the value of the expression. Round your result to two decimal places. (Enter your answer in radians.) arccot(−10)

zepdrix (zepdrix):

\(\Large\bf \color{royalblue}{\text{Welcome to OpenStudy! :)}}\)

OpenStudy (anonymous):

thank you!

zepdrix (zepdrix):

Hmm this question is kinda tricky :[ Lemme post my thoughts and see if I have the right idea here.

zepdrix (zepdrix):

\[\Large\bf\sf arccot(-10)\quad=\quad \theta\]\[\Large\bf\sf \cot \theta\quad=\quad -10\qquad\implies\qquad \tan \theta\quad=\quad -\frac{1}{10}\]\[\Large\bf\sf \arctan\left(-\frac{1}{10}\right)\quad=\quad \theta\]\[\Large\bf\sf \implies\qquad arccot(-10)\quad=\quad \arctan\left(-\frac{1}{10}\right)\] I did this because most calculators have an inverse tangent button, but not an inverse cotangent button. So now we can put it into the calculator.\[\Large\bf\sf \arctan\left(-\frac{1}{10}\right)\quad\approx\quad -0.099668652\]This measure is in radians. The output is in the 4th quadrant. No bueno. We are restricting our angle to the first 2 quadrants (0,π). Tangent and Cotangent are periodic in π so we can add π to our angle to get the correct output. \[\Large\bf\sf -0.099668652+\pi\quad\approx\quad 3.041924001~Radians\]Which is an angle between 0 and π.

zepdrix (zepdrix):

Grrr I hope I'm doing this right :p